OpenStudy (anonymous):

Solve the system, using substitution. y = 1 - 5x and 5x + y = 8 A. no solution B. infinitely many solutions

OpenStudy (anonymous):

Substitution means that we substitute a variable (In this case y) for an equation. Basically, we know 5x + y = 8, and we also know the value of y! The equation tells us that y = 1 - 5x. So we can substitute this in, like so. 5x + y = 8 5x + (1 - 5x) = 8. The brackets shouldn't go in for a real question. I'm just putting them in to show you exactly what's being substituted. In a real question, you should write it like this: 5x + 1 - 5x = 8. Now we have 5x and a -5x, so those add up to nothing. 1 = 8. Because 1 does not, in fact, equal 8, this question has no solutions.